First 2016 field update brought to you by Diana Wall

Greetings from the Crary.  Research is going full steam and the end of the season approaches, signaled by the arrival of the Coast Guard Ice Breaker. We had a string of summer days last week with spectacular weather conducive to getting our science done! The town is filling up with new arrivals from the field camps headed north and vessel offload personnel are soon coming to town. Tonight, Jeb and Byron will give the Sunday night lecture and we are looking forward to that. 


504-Glacier Team:  Adam has been at F6 and working on his hyporheic zone sampling and experiments. Anna has finished her work and has headed north. Forest is working with the stream team.

505-Limno team: Amy continues work on the Fryxell and Bonney moat bioassay experiment, and is processing DIC samples from the lakes, packing samples for shipment, and cleaning up from the season. The rest of the limno team headed north after a very successful season, and we miss them.

506- Stream team:  The stream team continues their stream monitoring and Algae Ops is in full swing with the arrival of Aneliya. Christa has completed her field season and is getting ready to head north.

507- The Wormherder team has been working at full force: the P3 was sampled and Byron, Jeb and Dan with some help from Adam, got the pulse/ press water treatments on the plots. The Biotic Effects Experiment, Stoichiometry Long term Monitoring experiment and Elevational Transects (LH, Garwood, Miers) have all been sampled and treatments applied where necessary. Scottnema data is looking very interesting thanks to Ashley, Tandra, Summer and Walter! We will make the vessel deadline this week, which in previous years has been a challenge due to weather delays. Some of the team (Jeb, Tandra) head north and will be missed (a lot). Ruth and Jess are busy with soil C02 on the Stoich plots and then will work on some Aeolian connectivity studies. Next week the BSNE’s, other sampling, plot maintenance and cargo continues.

509 – Geochem team:  Kathy, Elsa and Sydney are making progress on analytical work and completed packing of samples for the vessel deadline. Sydney and Elsa are back in town after helping Adam with his sampling and they also conducted a diurnal sampling for VonGuerard and Andersen Creek this week.

 Rae and Renee continue to provide support for all the teams from Lake Hoare Camp. They consistently provide plentiful and tasty food and a welcoming home and source of knowledge to the MCM and other science teams working in the valleys. 

 Good vibes continue to buzz through the labs. I think everyone is anxious and excited for the final push through the season. The weather was positive with no travel/flight cancelations over the weekday period. All MCM-LTERers appear dedicated to doing high quality work and getting the data and samples home. Seeing teams help each other to assure all LTER field science is completed before leaving, is a wonderful attribute of the MCM LTER.

We’ve had some friendly chats with Tom Wilch, who is the NSF science rep in town, and we are looking forward to chatting with Polly Penhale who arrived last night.


Bad news: 505 – Limno team: Amy attempted to deploy the West Lobe sediment trap, but was unsuccessful due to failure of the trap motor. The trap will be over-wintering in the vestibule of Lake Bonney camp.

UGLY - Nothing ugly here but the beards at the Mustache Roulette contest last night – the weather and science have been excellent!!

 All of us here send our regards and best wishes.

Article Section: